AO Commissioning on 14 October 2011

Observer(s):C. Arcidiacono, L. Busoni, A. Puglisi, JC. Guerra, K. Boutsia (LBTO)

Telescope Operator: D. Huerta (LBTO)

SW Support: T. Sargent (Tucson)

Telescope Support: M. Midkiff (LBTO)

Instrument Support: J. Morris (LBTO)

Summary

Clear sky when we open and winds <5m/s.
We had no particular problems with the telescope tonight and we made good progress with the AO tests
In particular the AO group made SW improvements, test dithering, test of ADI procedure, test off-axis aquisition, test of AO close loop on an extended source, focus of AO. For more details look the FLAOwiki and Lorenzo's email.
One standard star has been observed (BS9107) in open and closed loop in 3 filters (JHK).
Sky flats in JHK were obtained during sunset twilight and a set of darks for calculating corquad coefficients during dawn.

Details

00:20 We started taking some dome flats with the lights on - but while the scripts was executing the lights were turned off and we opened the dome for skyflats.
  • during the sequence we had to change filters because the dome is open.
  • the observe script was interrupted giving error :
.===> Error in pisces_send_request: Readu ERROR
        Either the network is down or the Pisces IRserver died
        You could also retry the command and a new connection will be attempted
  • this is probably due to the change of filters during the observe sequence.

00:36 starting sky flats
  • we did not manage to get narrow band (we started too late)
  • sky flats in Ks - starting with 8000counts (7sec) - images pisces.20111014.0086.fits - *.0097.fits
  • sky flats in H (1sec to test) - 13000counts (5sec) - images pisces.20111014.0104.fits - *.0109.fits
  • sky flats in Ks - starting with 5000counts (7sec) - images pisces.20111014.0110.fits - *.0115.fits
  • sky flats in H - 10000counts (5sec) - images pisces.20111014.0116.fits - *.0121.fits
  • sky flats in Ks - 3000counts (7sec) - images pisces.20111014.0122.fits - *.0127.fits
  • sky flats in H - 6000counts (5sec) - images pisces.20111014.0128.fits - *.0133.fits
  • sky flats in J - 14000counts (7sec) - images pisces.20111014.0135.fits - *.0140.fits
  • sky flats in H - 4000counts (5sec) - images pisces.20111014.0141.fits - *.0158.fits
  • sky flats in J - 4000counts (7sec) - images pisces.20111014.0159.fits - *.0171.fits

01:07 finished with skyflats but still too early to point a star for AO tests.
  • we close the dome and do some darks - images pisces.20111014.0172.fits - *.0220.fits

01:21 Open dome again to start observations.
  • checking pointing and collimation

02:00 preset to m15
  • preset failed because of error in the catalog ( Rmag was 0)
  • preset failed for ccd47 error
  • here they will test focus in different filters and do photometry tests
  • no star is found by the AO system

02:25 pisces server had to be restarted
  • the observe routine gave error with the filter: ====> Error in pisces_set_camera: Current filter 5 not found in list
  • we have disabled the filters in the observe routine. We will keep changing the filters outside the scripts.

02:30 after disabling the filter, scripts work and we go on with m15.
  • dimm - 1.0"

2:30 Preset to M15. The preset to M1 failed three times. The first one about darks ccd47, the second and the third about mistmatch of the magnitude expected from the requested.
  • Resend preset. Ok.
  • REquest an image to pisces the observe script reports an error. Error about pisces status. The problem came out because Pisces server is dead.
  • Restart the Pisces server and restart of the IDL, the next image went ok.
  • Applied dither pattern. ok.
  • Setting the procedure to determibne the focus of the AO system at differetn filters.

3:45 Resend the preset.101
  • need to solve the issue with the different filter focus positions in AO.
  • here the AO group has tried dither patterns in different filters, of the field with the star and a nearby sky field.
  • this set will be used also to test photometry
  • they have observed the same target in closed and open loop.
  • Close loop running without any problem (applying different dither scripts) for about 30 minutes.

04:35 preset to HR8799
  • here we will test ADI. Rotator stop
  • dimm 0.7''
  • During one set of images at the second image, the observe reported an error message: *Error in pisces sending request: Either the network is down or the pisces IRServer is died.... The same set was sent succesfull.

06:43 finished with ADI - stay on the same target to test the off-axis presets
  • there are problems with the off-axis presets
  • while they figure out how to slove them we do a standard

Standard BS9107 open and closed loop

08:54 Preset to standard BS9107.
  • we did dice5 in 3 filters (JHK)

09:18 Preset close loop BS9107. Failed.
  • Preset to AO system failed. Star not found.
===> Error in irc_runao
        err = Resmsg[0]  = RunAO result status: Error
        err = Resmsg[1]  = RETRY: No stars found in TV frame (-20001) WFSARB_STAR_NOTFOUND
        err = Resmsg[2]  = AcquireRef[437] failed - RETRY: No stars found in TV frame (-20001) WFSARB_STAR_NOTFOUND
  • preset failed again - star not found and I get error: ===> Active Optics loop is not running
  • the star was transiting... maybe that's why we see the target at the edge of the Pisces detector (x,y: 400,54)
  • send preset again - probably the guide star is too close and it is vignetting the AO guider... so we cancel the preset and try again with another GS.

09:41 Before starting the dither pattern, we realised that the AO internal focus had to be changed for the filter used.
  • at this moment using AO before every filter change we need to resend the preset after inserting the correct AO focus.
  • for the 3 filters in closed loop we will do: preset - dither - change filter - preset - dither.

10:11 repeat BS9107 in open loop
  • during the previous set the star was transiting and some images are elongated
  • we repeated all 3 filters with a larger dither pattern

Back to AO tests
10:29 Preset to NGC2273 (mag=15)
  • this set is to test guiding on an extended object.

10:58 preset failed. Active optics not ready. - need to correct pointing. Resend the preset and dither. The AO loop is close with the core of the galaxy. The autogain procedure is removed to avoid problems during the preset AO.

The problem found observing this object is that the observations request a sky, the sky is defined as a part of the sky without stars, so we can't build a catalogue to use with Observe routine. The work around is to point the telescope from the TO console and after we can run the Observe only with getimage and dither . We don't mind about the zero position of the telescope to start the dither pattern.

The system stopped due by secondary stack, the active optics corrections were removed to allow a new preset and collimation.

12:17 Resend the preset. Ok. Observations stop at 13:00

13:05 Closing Dome.

13:10 Set of long darks. Close dome and lights off.
  • 600sec darks still have high counts even though the lights are off. Is pisces getting warm?
  • pisces temp: 74.99 - this is not warm...

Comments on observe routines

  • Today we realised that images are saved twice because we run pisces_display_image. Actually the copy outside the *obs directory is saved in the directory from which the pisces_display_image has been launched.
  • While in open loop there is a pop-up window every few seconds asking if we want to go on with the script. Probably o pop-up window just once at the beginning of the script would be sufficient as warning.
  • In the image headers the IMAGETYPE info is not saved. It would also be useful to have the name of the object in the header.
  • During IR observation empty fields are necessary for sky subtraction. Since there are no stars in those fields it is not possible to create a catalog and do a preset using observe. What is the best strategy in this case?

-- KonstantinaBoutsia - 14 Oct 2011
Topic revision: r12 - 14 Oct 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback